OpenStudy (natsent06):

17% of a number is 153. What is 1% of the number?

OpenStudy (anonymous):

of a number... Which number?? Let us assume that number is \(x\).

OpenStudy (anonymous):

So : 17% of a number will be : \[\frac{17}{100} \times x\]

OpenStudy (anonymous):

17% is written as : \[17 \% = \frac{17}{100}\]

OpenStudy (anonymous):

"of" means multiplication ie \(\times\)

OpenStudy (anonymous):

and number we have let as : \(x\) : So, whole expression, 17% of number will be : \[\frac{17}{100} \times x\]

OpenStudy (anonymous):

Allergic to reply? or you don't have keyboard? or don't want to talk to me?

OpenStudy (natsent06):

i dont get it im confused

OpenStudy (anonymous):

I have not done anything yet, just see my replies above, and tell what are you not getting buddy? you can tell, we are friends here so don't hesitate. :)

OpenStudy (anonymous):

I am just decoding your question for you so that you can understand it line by line or word by word. :)

OpenStudy (natsent06):

should i give u my answer choices

OpenStudy (anonymous):

Well, it is so simple that I won't need any. :P Just kidding, I am no good at math, just learning it as we are talking or solving here. :) So, just follow the flow, we will get everything you require. :)

OpenStudy (anonymous):

Be honest, interact fully with me and I assure you that you will get it 100%, and if you don't get, then I will call some other friend of mine to help you. :P

OpenStudy (natsent06):

A) 26.01 B) 9 C) 900 D) 170

OpenStudy (anonymous):

Hey don't look at answer choices, just solve it step by step, we will get it for sure.

OpenStudy (anonymous):

So, let us start with, so read my sentences carefully with full concentration. :)

OpenStudy (anonymous):

your question says : `17% of a number is 153 ` I am decoding it word by word.

OpenStudy (anonymous):

As I said : `17%` means : \[\frac{17}{100}\]

OpenStudy (anonymous):

Suppose you are given : `15% of a number`, then `15%` is written as: \[\frac{15}{100}\]

OpenStudy (anonymous):

So, if you are getting that reply with : yes or yep or yup or something you feel good and if you are not getting it, then reply with : No, you are bad, you don't how to teach, or anything with which you can deny. :P

OpenStudy (natsent06):

Hokeyp

OpenStudy (anonymous):

What is this language? Which one? If I am giving freedom to you, then you will type anything? :P

OpenStudy (natsent06):

yush

OpenStudy (anonymous):

And after it says : `of a number` : So, `of` means : Multiplication sign means : \(\times \) yush or not?

OpenStudy (natsent06):

yush

OpenStudy (natsent06):

? @waterineyes

OpenStudy (xapproachesinfinity):

Hi buddy you still need help

OpenStudy (anonymous):

154 - 17%, use a calculator